# Fast Subspace Fluid Simulation with Temporal-Aware Basis

### v1

@@ 4.2. Learning from Noisy Data with Nonlinear Optimization
Fluid behaviour is inherently chaotic, particularly in highly turbulent systems. Additionally, discretization of such states leads to noisy data, particularly as structures approach the Nyquist limit of the sampling grid. Further, simulation speed and artist directability is of primary concern in graphics application above simulation accuracy, often leading to simulations with CFD-condition-violating large timesteps and early termination of iterative algorithms. This further leads to degraded simulation results. As such, we note that input fluid data in general will be highly noisy. Standard DMD, as shown in Equation 7, directly fits the data from the $i$-th frame to the ($i+1$)-th frame without considering the (potentially coherent) relationships between the $i$-th frame and the ($i+2$)-th or subsequent frames. This short time horizon makes it highly sensitive to noise, both spatially as the fluid state changes rapidly at the Nyquist limit, and at the operator level. As a result, standard DMD imposes high requirements on data quality.
To address this issue, we choose to use OptDMD (Askham and Kutz, 2018) for our graphics applications. OptDMD processes all snapshots simultaneously, significantly reducing the impact of noise by looking at the signal over a longer time horizon. By considering the $i$-th frame, the ($i+1$)-th frame, the ($i+2$)-th frame, and so on together, the influence of random noise is mitigated. To achieve this, OptDMD transforms the training of the reduced Koopman operator into an exponential data fitting problem and utilizes the variable projection method to solve this optimization problem. First, we perform an SVD of the dataset $\bm{X}$ to get $\bm{U_{r}},\bm{\Sigma_{r}},\bm{V_{r}}$ such that $\bm{X_{r}}=\bm{U_{r}}\bm{\Sigma_{r}}\bm{V_{r}^{T}}$. Then, we can construct the optimization problem as:
(12) $\operatorname*{arg\,min}_{\bm{\alpha},\bm{B}}\;\|\bm{\bar{V_{r}}\Sigma}-\bm{\Phi}(\bm{\alpha})\bm{B}\|_{F}$
where $\bm{\bar{V_{r}}}$ denotes the element-wise complex conjugate of $\bm{V_{r}}$, $\bm{\alpha}$ is an initial guess.
To solve this in reasonable runtime, we need a decent guess at $\bm{\alpha}$. To do so, we simply solve a single instance of the base DMD method (Equation 6) to find the eigenvalue matrix $\bm{\Lambda}$ and set $\bm{\alpha}=\bm{\Lambda}$.
The matrix $\bm{\Phi}$ can then be computed as:
(13) $\bm{\phi_{i}}=\frac{1}{\|\bm{U_{r}}\bm{B^{T}}(:,i)\|_{2}}\bm{U_{r}}\bm{B^{T}}(:,i)$
where $\bm{\phi_{i}}$ is the $i$-th colomn of matrix $\bm{\Phi}$.
@@ 4.3. Memory Overhead Optimization
DMD can effectively reconstruct datasets, but its training process requires a significant amount of memory due to the SVD. When the dataset is large, it becomes necessary to store an $N\times N$ matrix in memory, where $N$ represents the dimensionality of a snapshot, thus demonstrating a significant memory requirement.
We note that previous DMD literature either has much smaller datasets afforded by 2D data, or perform training on large clusters (Schmid, 2010; Proctor et al., 2016; Askham and Kutz, 2018; Sashidhar and Kutz, 2022). Uniquely in graphics, we expect algorithms to be able to simulate 3D examples (i.e. large $N$) but run on consumer hardware (i.e. small memory).
@@
Figure 4. Editing Temporal Dynamics of the Plume with Bunny with the Koopman Operator Approximation. In this experiment, we demonstrate the impact of changing the moduli of low-frequency and high-frequency modes in a 4:1 (left) and 1:4 (right) ratio. As mentioned in Section 6.1, in real applications, users can modify and manipulate the dynamic of different scales of vorticity by adjusting the reduced-order parameters.
[…]
4.4. Arbitary Time Step
@@
where $k$ is an integer of how many time steps we want to simulate forward. This can map any state at time $t$ to the state at time $t+k\Delta t$ in one matrix multiplication. This returns to the idea that the generator $\bm{\mathcal{K}}$ produces a family of operators $\{\bm{\mathcal{K}}_{t}\}$ that represent the flow map from $\bm{u}_{0}$ to $\bm{u}(t)$.
@@ 4.6. Encompassing External Forces in Reduced Space
where $\bm{B}_{i}$ and $\bm{q}_{i}$ can be constructed differently based on the settings of various scenarios. $\bm{f}_{j}$ $\in\mathbb{R}^{N\times 1}$ represents a user-modifiable external force. For example, in the 2D standalone plume scenario, we can define $\bm{q}_{1}\in\mathbb{R}^{M\times 1}$ as the density field, where $M$ represents the number of grid cells, and rewrite its impact on the velocity field in matrix form $\bm{B}_{1}\in\mathbb{R}^{N\times M}$. Similarly, we can define $\bm{q}_{2}\in\mathbb{R}^{M\times 1}$ as the temperature field, and rewrite its impact on the velocity field in matrix form $\bm{B}_{2}\in\mathbb{R}^{N\times M}$. These effects can then be projected onto the reduced space as $(\bm{\Phi}^{+}(\bm{B}_{1}\bm{q}_{1}+\bm{B}_{2}\bm{q}_{2}))$.
Figure 5. Comparison of PCA and Our Method for Low-Rank Flow Reconstruction. The leftmost column shows the reference high-resolution simulation, while the right grid presents reconstructions at different rank truncations ($r$). Each pair in the grid compares Principal Component Analysis (PCA) (left) and our method (right). Lower ranks ($r=2,9$) fail to capture large-scale turbulence while increasing $r$ improves accuracy. Our method retains more detailed structures at lower ranks than PCA, demonstrating improved efficiency in capturing complex flow dynamics.
[…]
@@ 5. Evaluation and Results
In this section, we demonstrate the results of our approach on baseline examples and provide detailed numerical evaluations. We conduct all numerical experiments on a Linux workstation equipped with a 2.10GHz 32-core Intel CPU featuring 120GB of RAM. We implement our method exclusively in Python, and our DMD implementation is a heavily optimized, memory-efficient version of PyDMD (Ichinaga et al., 2024). We use Taichi (Hu et al., 2019) for interactive examples and parallelization of DMD model inference, and NumPy (Harris et al., 2020) for numerical validation. Please refer to Table 2 for the experiment setup of the examples presented in this section.
@@ 5.2. Generalization With Vorticity Confinement
Figure 6. Comparison of Relative Error and Vorticity Confinement Between PCA (2013) and DMD (Ours). Left to right: (1) relative error over time for PCA (150 basis) (dashed) and Ours (50 basis) (solid), showing comparable or lower error with fewer basis functions. (2) reference high-resolution simulation with vorticity confinement $1.5$, with a zoomed-in region marked. (3) comparison of the zoomed region in (2) under novel unseen vorticity confinement force $1.51$, $1.6$, $2.5$.
[…]
@@ 5.5. Independence on Simulation Schemes
Our reduced simulation method is inherently simulator-agnostic, allowing it to work seamlessly with a variety of fluid solvers. This flexibility arises from the fact that our method models only the transitions between successive fluid states in the reduced space, rather than being tied to the specific equations or numerical schemes of a given solver. This allows us to apply our method to any fluid simulator without the need for additional adjustments. For instance, in our experiments, we used both the MacCormack (Selle et al., 2008) + Reflection (Zehnder et al., 2018) (MC+R) solver and a Lattice Boltzmann Method (Chen and Doolen, 1998) (LBM) solver. They are fundamentally different in their discretization and numerical operations. Specifically, the LBM solver does not directly solve the Navier-Stokes equations; instead, it solves the approximated Boltzmann equation, which models fluid dynamics at the mesoscopic scale using particle distribution functions. Our method overcame these challenges by relying solely on the data produced by the solver. In other words, our method is equation-free. We detailed the base simulators used for each example in Table 2, where our approach successfully reconstructs the velocity fields in all cases.
@@
Reversibility ($\triangleright$ Fig. 8) $512\times 512$ 2D Open $20$ MC
@@ 6.1. Direct Editing Temporal Dynamics
Figure 7. Editing temporal dynamics of Kármán Vortex Street with the Koopman Operator Approximation. The modifications are applied to the DMD basis coefficients: (a) Scaling the modulus of the DMD basis by factors of 0.5, 1.0, and 1.5, affecting overall amplitude; (b) Adjusting the real part of $\bm{\Omega}$, influencing growth and decay rates of modal contributions; (c) Modifying the imaginary part, altering phase dynamics and wave propagation characteristics.
[…]
Figure 8. Reversibility of Flows with Inversed DMD Operator. We compare the reconstruction of two distinct fluid flows using Dynamic Mode Decomposition (DMD). The top row in each panel shows the velocity L2-norm of the field used to train the DMD, while the second and third rows depict the temporal evolution of the reconstructed flow fields as applied to an initial density field. The forward-time training phase is followed by a backward-time testing phase to assess predictive accuracy when advecting backward in time. The bottom plots show the evolution of kinetic energy over time. From the buoyant case, we observe the inverted DMD operator $\bm{A^{-1}}$ can still reasonably trace backward in time without compromising much visual quality. The vortical case exhibits a more challenging example where the symmetry should be reconstructed backwards in time. We see that the inverse operator indeed recovers this symmetry, with some acceptable levels of incurred noise. Bottom plots show the evolution of the total kinetic energy over time, demonstrating that our inverse operator actually correctly reverses the arrow of time, reversing the dissipation-related entropy increase over time. Decreasing kinetic energy also validates the physical plausibility of our result.
[…]
Since our method approximates Equation 11 with a linear operator in the full space, this allows us to transform the operator acting on the velocity field into the evolution of different modes under a linear operator. Therefore, we can directly edit the temporal dynamics of the fluid system by modifying the modes of the reduced Koopman operator $\bm{\hat{K}}$: we set $t_{0}$ to be the initial time, $\bm{\Omega}=\nicefrac{{\log(\bm{\Lambda})}}{{\Delta t}}$, where $\Delta t$ is the time step of the dataset. With this, we can rewrite Equation 9 in the following form:
@@
As shown in Equation 16, our method decomposes a simulation sequence into modes with different growth/decay rates and frequencies. The growth/decay rate of a mode is reflected in $r_{i}$, where a larger $r_{i}$ indicates a higher growth rate (or a lower decay rate), and vice versa. The frequency of a mode is represented by the absolute value of $\theta_{i}$, with a larger absolute value corresponding to a higher frequency mode, and vice versa. Furthermore, the different modes are decoupled, allowing for the adjustment of the relative proportions between modes. As a result, these properties provide the artist with powerful tools to edit the simulation playback. The artist can modify the overall velocity field by adjusting the proportion ($w_{i}$), growth/decay rate ($r_{i}$), and frequency ($\theta_{i}$) of specific modes. In the experiments, we directly adjust the real part of $\bm{\Omega_{i}}$ to control $r_{i}$, modify the imaginary part of $\bm{\Omega_{i}}$ to control $\theta_{i}$, and vary the modulus of $\bm{\Phi_{i}}$ to control $w_{i}$.
@@ 6.2. Reversibility of the Reduced Simulation
Although physically-based fluid simulations have the capability to generate stunning visuals, when artists aim to direct the fluid’s evolution toward a predefined target shape, challenges arise. It is a long standing problem in the community that people aim to enable users with spatial control. In this example, we aim to enable users to do temporal control, motived by a prior work Oborn et al. (2018). Compared to previous work (2018) where the authors employ a self-attraction force to replace the arbitrary external forces, providing a stable, physics-motivated, but time-consuming approach, we propose a data-driven, fast, and easy to implement method to address the same problem.
@@
Similar to Equation 10, we could train the reduced Koopman operator on the forward simulation data and then apply the inverse reduced Koopman operator to reverse the simulation, given a state of the fluid system.
Figure 9. Upsampling and Generalization to Unseen Sequences with Trained DMD Operator. Two different input low-resolution fluid simulations (bunny and strawberry) are upscaled using the same DMD operator trained on a different velocity field. Initial velocity fields are seeded as moving down based on the input density field. Naive application of DMD shown in each middle column, and our augmented DMD upresolution method shown on the right columns. Schematic of our method presented on the far right. At each frame, we project the low-resolution artist-directed input into the low-order bases of our reduced representation, using these to replace the low-order terms of the DMD field. Notice that naive application of DMD simply moves towards the known input training data, while our augmented field matches the low-resolution input more closely, with extra high-order detail gained from the DMD operator.
[…]
@@
We experiment our approach on a simple buoyant flow setup (Fig. 8, left). Our dataset was initialized with a qian, a density field shaped like a round coin with a square hole, with the density value set to $1$. A density value of $1$ density field was driven by a velocity field where an upwards velocity of $0.3$ is set within the qian and downwards elsewhere. We run the simulation for $300$ frames to construct the dataset, and trained the DMD operator on this dataset. The inverse operator $\bm{\tilde{K}}^{-1}$ was then applied to the initial velocity field of the dataset at $t=0$ (frame 00). By iteratively applying the inverse operator, we obtained the velocity fields for the preceding frames, starting from frame $-1$, frame $-2$, and all the way back to frame $-300$. When examining the evolution of the density field from frame -300 to frame 300, it is evident that the velocity field remains consistently upward and smooth, indicating that our method is both reasonable and effective. Further analysis of the energy of the velocity field obtained through the inverse process and the velocity field from the dataset reveals a downward trend in energy, with a smooth and reasonable curve, consistent with fluids with dissipative properties. This demonstrates that our inverse operator has the ability to predict a physically-plausible velocity field prior to the dataset.

### v2

@@ 4.2. Learning from Noisy Data with Nonlinear Optimization
Fluid behaviour is inherently chaotic, particularly in highly turbulent systems. Additionally, discretization of such states leads to noisy data, particularly as structures approach the Nyquist limit of the sampling grid. Further, simulation speed and artist directability is of primary concern in graphics application above simulation accuracy, often leading to simulations with CFD-condition-violating large timesteps and early termination of iterative algorithms. This further leads to degraded simulation results. As such, we note that input fluid data in general will be highly noisy.
Standard DMD, as shown in Eq. (5), directly fits the data from the $i$-th frame to the ($i+1$)-th frame without considering the (potentially coherent) relationships between the $i$-th frame and the ($i+2$)-th or subsequent frames. This short time horizon makes it highly sensitive to noise, both spatially as the fluid state changes rapidly at the Nyquist limit, and at the operator level. As a result, standard DMD imposes high requirements on data quality.
To address this issue, we choose to use OptDMD (Askham and Kutz, 2018) for our graphics applications. OptDMD processes all snapshots simultaneously, significantly reducing the impact of noise by looking at the signal over a longer time horizon. By considering the $i$-th frame, the ($i+1$)-th frame, the ($i+2$)-th frame, and so on together, the influence of random noise is mitigated. To achieve this, OptDMD transforms the training of the reduced Koopman operator into an exponential data fitting problem and utilizes the variable projection method to solve this optimization problem. Similar to the procedure described in Sec. 3, we first perform Singular Value Decomposition (SVD) on the snapshot matrix $\bm{X}$ (see Eq. (4)) to obtain $\bm{U},\bm{\Sigma},\bm{V}$ such that $\bm{X}=\bm{U}\bm{\Sigma}\bm{V^{T}}$. Here, $\bm{\Sigma}$ is the diagonal singular value matrix, and $\bm{U},\bm{V}$ are orthonormal matrices whose columns are the left and right singular vectors, respectively.
With this decomposition, we can construct the optimization problem as:
(10) $\operatorname*{arg\,min}_{\bm{\alpha},\bm{B}}\;\|\bm{\bar{V}}\bm{\Sigma}-\bm{\Phi}(\bm{\alpha})\bm{B}\|_{F}$
where $\bm{\bar{V}}$ denotes the element-wise complex conjugate of $\bm{V}$, $\bm{\Phi}(\bm{\alpha})$ is the parameterized basis matrix with optimizable parameters $\bm{\alpha}$, and $\bm{B}$ is the coefficient matrix corresponding to the basis matrix. Here $\bm{\Phi}(\bm{\alpha})\in\mathbb{C}^{T\times r}$ is the Vandermonde matrix whose $(i,j)$-entry is $\bm{\Phi}(\bm{\alpha})]_{ij}=\alpha_{j}^{\,i-1}$ for $i=1,\dots,T$ and $j=1,\dots,r$; the vector $\bm{\alpha}=[\alpha_{1},\dots,\alpha_{r}]^{\top}$ collects the (discrete-time) DMD eigenvalues $\bm{\alpha}=\bm{\Lambda}$, initialized from the eigenvalue matrix $\bm{\Lambda}$ of a standard DMD solve (Eq. (4)).
Once the coefficient matrix $\bm{B}$ is estimated, we could compute each column of the basis matrix $\bm{\Phi}$ using:
(11) $\bm{\phi_{i}}=\frac{\bm{U_{r}}\bm{B^{T}}(:,i)}{\|\bm{U_{r}}\bm{B^{T}}(:,i)\|_{2}}$
where $\bm{\phi_{i}}$ denotes the $i$-th column of matrix $\bm{\Phi}$.
@@ 4.3. Memory Overhead Optimization
DMD can effectively reconstruct datasets, but its training process requires a significant amount of memory due to the SVD. When the dataset is large, it becomes necessary to store an $Nn\times Nn$ matrix in memory, where $Nn$ represents the dimensionality of a snapshot, thus demonstrating a significant memory requirement.
We note that previous DMD literature either has much smaller datasets afforded by 2D data, or perform training on large clusters (Schmid, 2010; Proctor et al., 2016; Askham and Kutz, 2018; Sashidhar and Kutz, 2022). Uniquely in graphics, we expect algorithms to be able to simulate 3D examples (i.e. large degrees of freedom) but run on consumer hardware (i.e. limited memory).
@@
Figure 4. Editing Temporal Dynamics of the Plume with Bunny with the Koopman Operator Approximation. In this experiment, we demonstrate the impact of changing the moduli of low-frequency and high-frequency modes in a 4:1 (left) and 1:4 (right) ratio. As mentioned in Sec. 6.1, in real applications, users can modify and manipulate the dynamic of different scales of vorticity by adjusting the reduced-order parameters.:
4.4. Arbitrary Time Step
@@
where $k$ is an integer of how many time steps we want to simulate forward. This can map any state at time $t$ to the state at time $t+k\Delta t$ in one matrix multiplication.
@@ 4.6. Encompassing External Forces in Reduced Space
where $\bm{B}_{i}$ and $\bm{q}_{i}$ can be constructed differently based on the settings of various scenarios. $\bm{f}_{j}$ $\in\mathbb{R}^{(Nn)\times 1}$ represents a user-modifiable external force. For example, in the 2D standalone plume scenario, we can define $\bm{q}_{1}\in\mathbb{R}^{M\times 1}$ as the density field, where $M$ represents the number of grid cells, and rewrite its impact on the velocity field in matrix form $\bm{B}_{1}\in\mathbb{R}^{(Nn)\times M}$. Similarly, we can define $\bm{q}_{2}\in\mathbb{R}^{M\times 1}$ as the temperature field, and rewrite its impact on the velocity field in matrix form $\bm{B}_{2}\in\mathbb{R}^{(Nn)\times M}$. These effects can then be projected onto the reduced space as $(\bm{\Phi}^{+}(\bm{B}_{1}\bm{q}_{1}+\bm{B}_{2}\bm{q}_{2}))$.
Figure 5. Comparison of PCA and Our Method for Low-Rank Flow Reconstruction. The leftmost column shows the reference high-resolution simulation, while the right grid presents reconstructions at different rank truncations ($r$). Each pair in the grid compares Principal Component Analysis (PCA) (left) and our method (right). Lower ranks ($r=2,9$) fail to capture large-scale turbulence while increasing $r$ improves accuracy. Our method retains more detailed structures at lower ranks than PCA, demonstrating improved efficiency in capturing complex flow dynamics.:
@@ 5. Evaluation and Results
In this section, we demonstrate the results of our approach on baseline examples and provide detailed numerical evaluations. We conduct all numerical experiments on a Linux workstation equipped with a 2.10GHz 32-core Intel CPU featuring 120GB of RAM. We implement our method exclusively in Python, and our DMD implementation is a heavily optimized, memory-efficient version of PyDMD (Ichinaga et al., 2024). We use Taichi (Hu et al., 2019) for interactive examples and parallelization of DMD model inference, and NumPy (Harris et al., 2020) for numerical validation. Please refer to Tab. 2 for the experiment setup of the examples presented in this section.
@@ 5.2. Generalization With Vorticity Confinement
Figure 6. Comparison of Relative Error and Vorticity Confinement Between PCA (2013) and DMD (Ours). Left to right: (1) relative error over time for PCA (150 basis) (dashed) and Ours (50 basis) (solid), showing comparable or lower error with fewer basis functions. (2) reference high-resolution simulation with vorticity confinement $1.5$, with a zoomed-in region marked. (3) comparison of the zoomed region in (2) under novel unseen vorticity confinement force $1.51$ (magenta), $1.6$ (blue), $2.5$ (dark green).:
@@ 5.5. Independence on Simulation Schemes
Our reduced simulation method is inherently simulator-agnostic, allowing it to work seamlessly with a variety of fluid solvers. This flexibility arises from the fact that our method models only the transitions between successive fluid states in the reduced space, rather than being tied to the specific equations or numerical schemes of a given solver. This allows us to apply our method to any fluid simulator without the need for additional adjustments. For instance, in our experiments, we used both the MacCormack (Selle et al., 2008) + Reflection (Zehnder et al., 2018) (MC+R) solver and a Lattice Boltzmann Method (Chen and Doolen, 1998) (LBM) solver. They are fundamentally different in their discretization and numerical operations. Specifically, the LBM solver does not directly solve the Navier-Stokes equations; instead, it solves the approximated Boltzmann equation, which models fluid dynamics at the mesoscopic scale using particle distribution functions. Our method overcame these challenges by relying solely on the data produced by the solver. In other words, our method is equation-free. We detailed the base simulators used for each example in Tab. 2, where our approach successfully reconstructs the velocity fields in all cases.
@@
Reversibility ($\triangleright$ Fig. 8) $512\times 512$ 2D Open 2020 MC
@@ 6.1. Direct Editing Temporal Dynamics
Figure 7. Editing temporal dynamics of Kármán Vortex Street with the Koopman Operator Approximation. The modifications are applied to the DMD basis coefficients: (a) Scaling the modulus of the DMD basis by factors of 0.5, 1.0, and 1.5, affecting overall amplitude; (b) Adjusting the real part of $\bm{\Omega}$, influencing growth and decay rates of modal contributions; (c) Modifying the imaginary part, altering phase dynamics and wave propagation characteristics. :
Figure 8. Reversibility of Flows with Inversed DMD Operator. We compare the reconstruction of two distinct fluid flows using Dynamic Mode Decomposition (DMD). The top row in each panel shows the velocity L2-norm of the field used to train the DMD, while the second and third rows depict the temporal evolution of the reconstructed flow fields as applied to an initial density field. The forward-time training phase is followed by a backward-time testing phase to assess predictive accuracy when advecting backward in time. The bottom plots show the evolution of kinetic energy over time. From the buoyant case, we observe the inverted DMD operator $\bm{A^{-1}}$ can still reasonably trace backward in time without compromising much visual quality. The vortical case exhibits a more challenging example where the symmetry should be reconstructed backwards in time. We see that the inverse operator indeed recovers this symmetry, with some acceptable levels of incurred noise. Bottom plots show the evolution of the total kinetic energy over time, demonstrating that our inverse operator actually correctly reverses the arrow of time, reversing the dissipation-related entropy increase over time. Decreasing kinetic energy also validates the physical plausibility of our result.:
Since our method approximates Eq. (9) with a linear operator in the full space, this allows us to transform the operator acting on the velocity field into the evolution of different modes under a linear operator. Therefore, we can directly edit the temporal dynamics of the fluid system by modifying the modes of the reduced Koopman operator $\bm{\hat{K}}$: we set $t_{0}$ to be the initial time, $\bm{\Omega}=\nicefrac{{\log(\bm{\Lambda})}}{{\Delta t}}$, where $\Delta t$ is the time step of the dataset. With this, we can rewrite Eq. (7) in the following form:
@@
As shown in Eq. (14), our method decomposes a simulation sequence into modes with different growth/decay rates and frequencies. The growth/decay rate of a mode is reflected in $r_{i}$, where a larger $r_{i}$ indicates a higher growth rate (or a lower decay rate), and vice versa. The frequency of a mode is represented by the absolute value of $\theta_{i}$, with a larger absolute value corresponding to a higher frequency mode, and vice versa. Furthermore, the different modes are decoupled, allowing for the adjustment of the relative proportions between modes. As a result, these properties provide the artist with powerful tools to edit the simulation playback. The artist can modify the overall velocity field by adjusting the proportion ($w_{i}$), growth/decay rate ($r_{i}$), and frequency ($\theta_{i}$) of specific modes. In the experiments, we directly adjust the real part of $\bm{\Omega_{i}}$ to control $r_{i}$, modify the imaginary part of $\bm{\Omega_{i}}$ to control $\theta_{i}$, and vary the modulus of $\bm{\Phi_{i}}$ to control $w_{i}$.
@@ 6.2. Reversibility of the Reduced Simulation
Although physically-based fluid simulations have the capability to generate stunning visuals, when artists aim to direct the fluid’s evolution toward a predefined target shape, challenges arise. It is a long standing problem in the community that people aim to enable users with spatial control. In this example, we aim to enable users to do temporal control, motivated by a prior work Oborn et al. (2018). Compared to previous work (2018) where the authors employ a self-attraction force to replace the arbitrary external forces, providing a stable, physics-motivated, but time-consuming approach, we propose a data-driven, fast, and easy to implement method to address the same problem.
@@
Similar to Eq. (8), we could train the reduced Koopman operator on the forward simulation data and then apply the inverse reduced Koopman operator to reverse the simulation, given a state of the fluid system.
Figure 9. Upsampling and Generalization to Unseen Sequences with Trained DMD Operator. Two different input low-resolution fluid simulations (bunny and SIGGRAPH logo) are upscaled using the same DMD operator trained on a different velocity field. Initial velocity fields are seeded as moving down based on the input density field. Naive application of DMD shown in each middle column, and our augmented DMD upresolution method shown on the right columns. Schematic of our method presented on the far right. At each frame, we project the low-resolution artist-directed input into the low-order bases of our reduced representation, using these to replace the low-order terms of the DMD field. Notice that naive application of DMD simply moves towards the known input training data, while our augmented field matches the low-resolution input more closely, with extra high-order detail gained from the DMD operator.:
@@
We experiment our approach on a simple buoyant flow setup (Fig. 8, left). Our dataset was initialized with a density field shaped like the SIGGRAPH logo, with the density value set to 1. A density value of $1$ density field was driven by a velocity field where an upwards velocity of $0.3$ is set within the SIGGRAPH logo and downwards elsewhere. We run the simulation for $300$ frames to construct the dataset, and trained the DMD operator on this dataset. The inverse operator $\bm{\tilde{K}}^{-1}$ was then applied to the initial velocity field of the dataset at $t=0$ (frame 0). By iteratively applying the inverse operator, we obtained the velocity fields for the preceding frames, starting from frame 1-1, frame 2-2, and all the way back to frame 300-300. When examining the evolution of the density field from frame -300 to frame 300, it is evident that the velocity field remains consistently upward and smooth, indicating that our method is both reasonable and effective. Further analysis of the energy of the velocity field obtained through the inverse process and the velocity field from the dataset reveals a downward trend in energy, with a smooth and reasonable curve, consistent with fluids with dissipative properties. This demonstrates that our inverse operator has the ability to predict a physically-plausible velocity field prior to the dataset.